OpenStudy (anonymous):

. A truck traveling at an average rate of 45 miles per hour leaves a rest stop. Fifteen minutes later a car traveling at an average rate of 60 miles per hour leaves the same rest stop traveling the same route. How long will it take for the car to catch up with the truck? 45 minutes 3 hours 15 minutes 1 hour 15 minutes

OpenStudy (error1603):

suppose that Vt is average rate of truck and x its the time, while Vc is for the car and with its time be (x-1/4). make an equation according this information: Vt * x = Vx * (x-1/4) (because distance for both be same) 45x = 60(x-1/4) now, solve for x.... what u get

OpenStudy (error1603):

Hello

OpenStudy (anonymous):

x=1?

OpenStudy (error1603):

45x = 60(x-1/4) 45x = 60x - 15 right?

OpenStudy (error1603):

I have used distributive property to simplify for right side so, 60 multiplied by x = 60x and 60 multiplied by 1/4 = 15 right?

OpenStudy (mrnood):

The car is catching the truck at 15 mph How far ahead is the truck when the car starts? (hint it travels for 15 mins @ 45 mph) How long does it take the car to close the gap when catching at 15mph?

OpenStudy (anonymous):

now im completely confused.

OpenStudy (error1603):

i just give a hint. after u solve for x, then to get the time for the car to catch up with the truck, just subtracted by 15 minutes (because the car was late 15 minutes to the truck). u will get the answer.

OpenStudy (error1603):

Got it?

OpenStudy (error1603):

Not talking?

OpenStudy (error1603):

oooookay